# Bridging the Linguistic Divide: A Survey on Leveraging Large Language Models for Machine Translation

### v2

@@ 5 LLMs vs Encoder-Decoder across Languages
This section explores how large language models (LLMs) perform across languages and how they compare to dedicated machine translation (MT) systems. An early study found that ChatGPT lags behind traditional MT systems such as Google Translate, DeepL, and Tencent, but observed that GPT-4 performs competitively with these systems in zero-shot settings [34]. This indicates that some LLMs achieve performance on par with traditional MT systems for high-resource languages.
@@
When translating from Spanish to 11 indigenous South American languages, fine-tuned M2M100 was shown to consistently outperform GPT-4 [68]. The performance gap further increases when M2M100 is combined with techniques like kNN-MT [69]. The contrast between high-resource and low-resource languages is also highlighted by a multitask evaluation, where ChatGPT’s performance, although slightly lower, remains close to fine-tuned state-of-the-art models [70]. For instance, ChatGPT achieves 58.64 chrF++ compared to 63.5 for fine-tuned models in XXEn translation, and 51.12 chrF++ versus 54.4 for EnXX [71]. However, for low-resource languages, the difference becomes much more pronounced, with ChatGPT scoring 27.79 chrF++ compared to 54.9 for fine-tuned models in XXEn, and 21.57 versus 41.9 chrF++ for EnXX [53].
@@ 6 LLM based Evaluation
The performance of large language models (LLMs) on the MQM-2020 dataset has been evaluated by comparing their Spearman correlations with traditional machine translation (MT) metrics, such as ROUGE, BERTScore, and BARTScore [72]. The findings indicate that, when provided with appropriate instructions and demonstrations, LLMs can outperform these conventional metrics. However, the comparison is restricted to high-resource language pairs (English and Chinese), and more advanced metrics, such as COMET, which exhibit higher correlation with human judgments, were not included.
@@
Building on these approaches, Error Analysis Prompting has been introduced to combine error analysis with chain-of-thought reasoning for MT evaluation using LLMs [76]. Experiments reveal that, while LLMs achieve higher system-level accuracy, they fall short of traditional metrics, including embedding-based approaches such as BLEURT, COMET, and MetricX, at the segment level.
The distinction between system-level and segment-level evaluation is crucial in MT assessment. System-level evaluation refers to the overall performance of an MT system, typically measured by aggregating scores across multiple sentences or documents. In contrast, segment-level evaluation assesses the quality of individual translated sentences, offering a more fine-grained analysis. While LLMs demonstrate strong system-level accuracy, their lower segment-level accuracy suggests potential inconsistencies when evaluating individual translations, highlighting a key limitation compared to traditional metrics.
@@ 7 Summary and Recommendation
This survey has provided a comprehensive analysis of how Large Language Models (LLMs) are reshaping the landscape of machine translation (MT). The review highlighted the effectiveness of LLMs in generating synthetic data, leveraging few-shot and in-context learning, and fine-tuning through parameter-efficient techniques such as LoRA and QLoRA. While LLMs show competitive performance in high-resource language pairs, they still lag behind traditional MT systems for low-resource languages due to data scarcity and linguistic diversity challenges. Additionally, the evaluation of LLM-based MT demonstrates that although LLMs can achieve competitive system-level accuracy, they face limitations in segment-level consistency compared to traditional metrics like COMET.

### v3

@@ 5 LLMs and Encoder-Decoder across Languages
This section explores how LLMs perform across languages and how they compare to dedicated MT systems. An early study found that ChatGPT lags behind traditional MT systems such as Google Translate, DeepL, and Tencent, but observed that GPT-4 performs competitively with these systems in zero-shot settings [48]. This indicates that some LLMs achieve performance on par or even better with traditional MT systems for high-resource languages.
@@
When translating from Spanish to 11 indigenous South American languages, fine-tuned M2M100 was shown to consistently outperform GPT-4 [96]. The performance gap further increases when M2M100 is combined with techniques like kNN-MT [97]. The contrast between high-resource and low-resource languages is also highlighted by a multitask evaluation, where ChatGPT’s performance, although slightly lower, remains close to fine-tuned state-of-the-art models [98]. For instance, ChatGPT achieves 58.64 chrF++ compared to 63.5 for fine-tuned models in XXEn translation, and 51.12 chrF++ versus 54.4 for EnXX [44]. However, for low-resource languages, the difference becomes much more pronounced, with ChatGPT scoring 27.79 chrF++ compared to 54.9 for fine-tuned models in XXEn, and 21.57 versus 41.9 chrF++ for EnXX [69]. A similar study [99] identifies key bottlenecks, such as the asymmetry in translation quality between high-resource to low-resource directions and the poor generative capability of smaller models. The study finds that LLMs generally underperform for low-resource languages compared to traditional encoder-decoder NMT systems, which is also confirmed by [100] earlier. Further, they observe that generating En XX is more difficult compare to generating XXEn.
Given the poor state of performance in low-resource languages, the authors in [101] examine whether LLMs can learn to translate extremely low-resource languages using only grammar books. They split grammar books into parallel examples and grammatical explanations, then tested LLMs and a fine-tuned NMT model on translation, grammatical judgment, and gloss prediction tasks. They also introduced typological prompts summarizing high-level grammatical features. Their findings show that translation improvements come almost entirely from parallel examples, not grammatical explanations. Fine-tuned models performed as well or better than LLMs using grammar books. However, for linguistic tasks like grammaticality judgment and gloss prediction, typological prompts combined with parallel examples led to strong performance. The study concludes that for translation, collecting parallel data is more effective than relying on descriptive grammar, while structured grammatical knowledge is better suited for non-translation tasks. The authors in [102] investigates the persistent issue of the Curse of Multilinguality (CoM) in LLM-based multilingual MT, where scaling to more languages introduces performance bottlenecks. The authors identify a key phenomenon: asymmetry in linguistic conflicts and synergies—with translation from other languages to English (XXEn) suffering from conflicts, while English to other languages (EnXX) benefits from synergy. These issues primarily arise during the post-training phase rather than during multilingual pretraining. To address this, they propose a Direction-Aware Training (DAT) approach: applying separate training for XXEn directions to reduce conflicts and group multilingual training for EnXX to leverage synergy. They further introduce group-wise model merging (DATM) to improve scalability by reducing the number of LoRA adapters while maintaining translation quality. This work highlights the importance of targeted post-training strategies and provides a path toward building high-quality, cost-effective MT systems.
Towards building efficient LLM-based MT systems, a pre-trained decoder-only LLM can serve as the encoder, paired with a lightweight NMT-style decoder [103]. To enable compatibility between the LLM’s output and the decoder’s input, the intermediate representations from multiple LLM layers are integrated, transformed to a lower dimensional space, and optionally enhanced with additional encoder layers to capture bidirectional context. Training proceeds in two phases: initially, only the transformation layers and decoder are trained while keeping the LLM frozen; subsequently, the entire model is fine-tuned on ComMT, a comprehensive benchmark covering general, domain-specific, document-level, terminology-constrained, and post-editing translation tasks. This approach, called LaMaTE, demonstrates strong performance and generalization, while significantly improving efficiency—achieving up to 6.5× faster decoding and 75% reduction in memory usage—highlighting the potential of LLMs as encoders in NMT architectures.
A key distinction between LLMs and encoder-decoder models is their ability to generalize beyond sentence-level translation, offering rich contextualization, stylistic control, and flexibility via prompting. A recent study [79] reveals that this distinction erodes when LLMs are fine-tuned on parallel data. In many respects, fine-tuned LLMs begin to resemble traditional NMT systems: they lose control over output formality, struggle with context-sensitive disambiguation, and fail to generalize effectively to unseen domains without explicit retraining. While idiomatic translation often improves—particularly when initial performance is weak—other capabilities, such as formality control and document-level contextualization, suffer significant degradation. This degradation is not readily apparent when relying solely on standard evaluation metrics. For instance, COMET scores may remain stable or even improve while the model’s ability to follow stylistic constraints or disambiguate based on discourse context declines. This indicates a mismatch between what current MT metrics measure and what users value in real-world translation tasks. This convergence suggests that careless adaptation of LLMs risks negating their primary benefits. While LLMs initially surpass NMT systems in their versatility, indiscriminate fine-tuning can narrow their capabilities to those of more brittle translation architectures. This finding reinforces the need to develop fine-tuning strategies that preserve the emergent capabilities of LLMs rather than inadvertently discarding them.
@@ 6 LLM based Evaluation
The performance of LLMs) on the MQM-2020 dataset has been evaluated by comparing their Spearman correlations with traditional MT metrics, such as ROUGE, BERTScore, and BARTScore [104]. The findings indicate that, when provided with appropriate instructions and demonstrations, LLMs can outperform these conventional metrics. However, the comparison is restricted to high-resource language pairs (English and Chinese), and more advanced metrics, such as COMET, which exhibit higher correlation with human judgments, were not included.
@@
Building on these approaches, Error Analysis Prompting has been introduced to combine error analysis with CoT reasoning for MT evaluation using LLMs [108]. Experiments reveal that, while LLMs achieve higher system-level accuracy, they fall short of traditional metrics, including embedding-based approaches such as BLEURT, COMET, and MetricX, at the segment level.
The authors in [111] investigates the effectiveness of LLM in evaluating MT quality without requiring fine-tuning. The authors systematically examine what translation-related information, such as the source sentence, reference translation, identified errors, and annotation guidelines is necessary for LLMs to reliably assess MT quality. The findings highlight that the inclusion of both the source and reference translations is crucial for improving evaluation accuracy, while error annotations and detailed guidelines offer limited additional benefit. Larger models tend to benefit more from CoT prompting, though smaller models also perform competitively. However, LLMs are inconsistent in generating numerical scores, raising concerns about their reliability for automated MT evaluation. For a more robust evaluation, M-MAD [112] use a three-stage framework for via structured multi-agent collaboration. First, it decouples the MQM evaluation guideline into four distinct dimensions - accuracy, fluency, style, and terminology — allowing dimension-specific LLM agents to identify and annotate errors independently. In the second stage, each dimension undergoes a focused two-agent debate, where one agent supports and the other critiques the initial annotation across multiple rounds, fostering more accurate and balanced error assessments. Finally, a judge agent synthesizes the debated viewpoints and computes the overall MQM score using weighted counts of major and minor errors. This structured approach significantly enhances system-level correlation; however, non-LLM metrics fared better in segment-level.
@@ 7 Summary and Recommendation
This survey has provided a comprehensive analysis of how LLMs are reshaping the landscape of MT. It highlighted the effectiveness of LLMs in generating synthetic data, leveraging few-shot and in-context learning, and enhancing performance through parameter-efficient fine-tuning techniques such as LoRA and QLoRA. While LLMs demonstrate competitive results in high-resource language pairs, they continue to underperform compared to traditional MT systems in low-resource settings due to persistent challenges like data scarcity and linguistic diversity. Furthermore, the evaluation of LLM-based MT reveals that despite achieving strong system-level accuracy, LLMs often struggle with segment-level consistency, as measured by metrics such as COMET. Despite the progress, a number of open challenges persist, motivating future research directions as outlined below:
@@
It is important to emphasize that MT with LLMs is a vast and rapidly evolving area of research. Each aspect—be it data generation, learning paradigms, fine-tuning strategies, or evaluation—warrants a dedicated survey of its own. Due to space and scope limitations, this review focuses primarily on sentence-level translation. Consequently, while we aimed to cover a representative subset of the literature, some relevant works may not have been included.